# Regge trajectories and bridges between them in integrable AdS/CFT

### v2

@@ 2.1.4 Asymptotics and quantum numbers
$\mathbf{P}_{ab}\sim\#\,u^{\mathcal{N}_{a}+\mathcal{N}_{b}}\quad\text{as}\quad u\gg 1\,,$ (30)
@@
${\bf Q}_{ij}\sim\#\,u^{\hat{\mathcal{N}}_{i}+\hat{\mathcal{N}}_{j}-1}\quad\text{as}\quad u\gg 1,$ (32)
@@
$\hat{M}_{2}=\Delta_{\text{top}}-S_{\text{top}}+1\,,\quad\hat{M}_{1}=\Delta_{\text{top}}+S_{\text{top}}+2,$
where $r_{1}\equiv 1/2\,(p_{1}+p_{2})+q$, $r_{2}\equiv 1/2\,(p_{1}+p_{2})$, $r_{3}=1/2\,(p_{2}-p_{1})$ are the corresponding $SO(6)$ weights, and in fact the condition (34) translates into the unitarity bound satisfied by physical multiplets
@@
We notice that is the bottom component that gives the leading Regge trajectory in the sector we will consider. In the following, we will always plot only one representative for each multiplet, specifying which.
@@ 2.2 Weyl symmetries in the QSC
They appear as a symmetry of the QSC under simple Q-system transformations that multiply the “fermionic” indices of Q-functions, namely the indices of type $i,j=1,\dots,4$, by multiplication with one of the following matrices (each one defining a new solution):
$\left(\mathbb{M}_{1}\right)_{i}^{\;j}=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&0&1&0\\ 0&-1&0&0\\ 0&0&0&1\end{array}\right),\;\;\;\;\text{or}\;\;\;\;\left(\mathbb{M}_{2}\right)_{i}^{\;j}=\left(\begin{array}[]{cccc}0&0&0&1\\ 0&1&0&0\\ 0&0&1&0\\ -1&0&0&0\end{array}\right),$ (41)
where upper indices should be transformed by multiplying by the inverse transposed of these matrices. From the QQ relations point of view, these preserve the form of the Q-system [^14] In particular, this follows from the fact that $\mathbb{M}\cdot\kappa\cdot\mathbb{M}^{T}=\kappa$., and at the same time preserve the analytic properties with an appropriate redefinition of the gluing matrix $\mathcal{G}\rightarrow\mathbb{M}\cdot\mathcal{G}\cdot\mathbb{M}^{-1}$. However, these transformations also correspond to a change in asymptotics and thus generate a new solution.
In particular, notice that conjugation by $\mathbb{M}_{1}$ corresponds to swapping the indices [^15] A part for the appropriate ±$\pm 1$ signs coming from the components of $\mathbb{M}_{i}$. These signs are needed to preserve the QQ relations, but play no role in the present argument on the redefinition of the charges.
$i=2\,\longleftrightarrow\,i=3,$ (42)
and thus the charges of the new solution are obtained from the ones of the original ones by the swap (the notation will be clear in a minute):
$\texttt{W}_{S}:\;\;\;\hat{M}_{1}\,\longleftrightarrow\,\hat{M}_{2}.$ (43)
@@
$\texttt{W}_{\Delta}:\;\;\;\hat{M}_{1}\,\longleftrightarrow\,-\hat{M}_{2}.$ (45)
@@
$\texttt{W}_{\Delta}:\;\;\;\;\left(\Delta_{\text{bottom}},S_{\text{bottom}}\right)\,\longrightarrow\,\left(3-\Delta_{\text{bottom}},S_{\text{bottom}}\right),$ (46)
@@
$\texttt{W}_{S}:\;\;\;\;\left(\Delta_{\text{top}},S_{\text{top}}\right)\,\longrightarrow\,\left(\Delta_{\text{top}},-1-S_{\text{top}}\right).$ (47)
These are individually Weyl reflections that leave the conformal algebra invariant. [^16] In general dimension, the two Weyl reflections would act us $\texttt{W}_{\Delta}:\;\Delta\,\longleftrightarrow\,d-\Delta$, and $\texttt{W}_{S}:\;S\,\longleftrightarrow\,2-d-S$. Following the CFT literature, we will call the two transformations the shadow transform and the spin-shadow. The reason why one acts on the bottom component and the other on the top component is due to the supersymmetric structure. Indeed, these two particular transformations, acting at the levels of the multiplets specified in (46) and (47), are Weyl reflections of the whole $OSp(6|4)$ superalgebra, as can be seen from the form of the Casimir eigenvalue, see e.g. Binder:2021cnk .
@@
$\tau=\left.\Delta-S\right|_{\text{top}}=\Delta_{\text{top}}-S_{\text{top}},\;\;\;\tilde{\tau}=\left.\Delta-(\texttt{W}_{S}[S])\right|_{\text{top}}=\Delta_{\text{top}}+1+S_{\text{top}}.$ (48)
@@ 3.1 A symmetric sector
This symmetry then percolates to other elements of the construction. In particular, by studying (7), we realize that now $Q^{a|k}\kappa_{ki}$ and $\kappa^{ab}Q_{b|i}$ satisfy the same difference equation. Identifying solutions via their asymptotics, one can choose an appropriate normalization of the Q functions such that (see, Bombardelli:2017vhk )  [^17] We are not adopting exactly the same conventions as in Bombardelli:2017vhk . The canonical normalization of Q-functions used here for the LR-symmetric case differs from that paper. In the LR symmetric case we have $e^{\mathrm{i}\mathcal{P}}=\mathbf{s}=\pm 1$. A redefinition related to this sign connects the convention of this paper to the one in Bombardelli:2017vhk , i.e. $Q_{a|i}^{\text{here}}$ is equivalent to $Q_{a|i}$ in Bombardelli:2017vhk if $\mathbf{s}=1$, while if $\mathbf{s}=-1$ we multiply the $a$ index by a diagonal matrix $\left\{\mathrm{i},-\mathrm{i},-\mathrm{i},\mathrm{i}\right\}$. :
@@
A second $\mathbb{Z}_{2}$ symmetry of the system is an internal symmetry acting on the spin chain representation of the states by a reflection across the midpoint of the spin chain. We will restrict our analysis to states that are eigenvalues of this transformation. At the level of Q functions this $\mathbb{Z}_{2}$ parity is realized very simply as the change of spectral parameter $u\leftrightarrow-u$ of the $\mathbf{P}$ functions. [^18] We should specify that the transformation $u\leftrightarrow-u$ should be performed on a sheet where the branch cut on the real axis is drawn as short, i.e. along $[-2h,2h]$.
For parity-symmetric states, this sends the QSC solution to itself which means that the set of $\mathbf{P}$ functions is invariant under $u\leftrightarrow-u$ up to rescalings. [^19] Notice also that this symmetry implies that for parity-symmetric states the large-$u$ expansions of all quantities will go in steps of $1/u^{2}$ only. In practice, since they are defined as series in $1/x(u)$ and $x(-u)=-x(u)$, this means that they just pick a sign, based on the charges in their asymptotics, e.g. $\mathbf{P}_{12}(-u)=(-1)^{M_{1}}\mathbf{P}_{12}(u)$ since the asymptotics is $\sim u^{-M_{1}}$ and $M_{1}\in\mathbb{Z}$. We can represent these changes of sign conveniently as a Q-system transformation, i.e.
@@
Notice $\mathbb{R}\kappa\mathbb{R}^{T}=\kappa$, which ensures that this symmetry is compatible with LR symmetry.
In presence of spin chain parity, it is very simple to construct the Q-functions satisfying the Q-system holomorphic in the lower half plane, i.e. the objects we already encountered, $Q_{a|i}^{\uparrow}$ and $\mathbf{Q}_{ij}^{\uparrow}$. Given the symmetry, they can in fact now simply be taken as
$Q_{a|i}^{\uparrow}=(\mathbb{R}^{-1})_{a}^{\,c}\;Q_{c|i}^{\downarrow}(-u),$ (59)
@@ 3.1.1 Constraints on the gluing matrix
Let us now see what additional conditions should be satisfied by the gluing matrix in presence of the two symmetries above. We start by rewriting the gluing equation (17), using the $u\leftrightarrow-u$ symmetry, as
@@
  • Two new constraints for the matrix $\Theta$ in this sector (see Appendix C), namely
@@ 3.2.1 Strategy of the derivation
Since $\mathcal{G}(u)$ is a periodic matrix of holomorphic functions with no singularities anywhere for finite $u$, it is severely constrained. In particular, if we require that it has non-growing asymptotics at $u\rightarrow\pm\infty$, then it should be a constant by Liouville’s theorem: this is the case of local operators. To have a nontrivial dependence on $u$, the most general class of functions which do not grow worse than exponentials are polynomials in $e^{\pm\pi u}$. We will make the following
@@
    $f(u)\equiv\mathcal{G}(u)\cdot\Theta\left(u-\mathrm{i}/2\right),$
    and impose that it can be parametrized in terms of 4 $\tau$-functions in the correct way. This equation should be studied asymptotically at both +$\infty$+ ∞ and $\infty$- ∞, using an appropriate expansion of $\Theta$.
  • Once the parameter space is sufficiently tiny, impose the quadratic constraint (65) to further increase the constraint range.
@@ 3.2.2 Review of the result for local operators
$\tau_{2,3}\sim 0,$ (75)
$\tau_{1,4}\sim\texttt{const}\,,$ (76)
@@
Using the strategy described above and in particular the algorithm explained in Appendix D, one gets as a result a matrix depending on the charges of the state (here given in $sl(2)$ grading) and just one unfixed parameter Bombardelli:2017vhk
$\mathcal{G}_{i\,\text{local}}^{\,\,\,j}=\begin{pmatrix}\mathrm{i}\sec\left(\pi\Delta_{sl(2)}\right)&0&0&\mathrm{i}\,\delta_{1}\\ 0&\mp\,\mathrm{i}&0&0\\ 0&0&\pm\,\mathrm{i}&0\\ -\mathrm{i}\,\delta_{2}&0&0&-\mathrm{i}\sec\left(\pi\Delta_{sl(2)}\right)\\ \end{pmatrix},$ (77)
@@
$-\mathrm{i}\,\text{e}^{-\imath\pi S_{sl(2)}}=\mp\mathrm{i},$ (78)
where this is the same $\pm$ sign appearing in (77). We see that the cases of even/odd-spin operators are explicitly disjoint.
As mentioned in Subsection 2.3, the above result is very general and is not relegated to the particular sector considered. Rather it is a consequence of the hypothesis of a gluing matrix independent of $u$. See Appendix D for a detailed discussion.
@@ 3.2.3 Result for non-local operators
To avoid spin quantization, we need to relax the asymptotic behaviour of $\tau$-functions (75)-(76), which will allow to make the gluing matrix dependent on $u$. Following the strategy described above, we focus on the simplest non-trivial assumption compatible with the enhanced periodicity of the sector (55), i.e.
$\tau_{2,3}\sim A_{2,3}^{\pm}\text{e}^{\pi\,|u|}+B_{2,3}^{\pm}\text{e}^{-\pi\,|u|},$ (79)
$\tau_{1,4}\sim C_{1,4}^{\pm}+D_{2,3}^{\pm}\text{e}^{-2\pi\,|u|}$ (80)
as $u\rightarrow\pm\infty$, where $A,B,C,D\in\mathbb{C}$ are some parameters. In this work we analyze in detail two disjoint configurations of the general ansatz (79), e.g.
@@
Notice $A_{3}^{\pm}=A_{2}^{\pm}=0$ reduces to what discussed in Section 3.2.2.
@@
Following the strategy discussed in 3.2.1, let us examine in detail the steps we used to fix the form of the $\mathcal{G}(u)$ in the case (81).
@@
    $\left(\begin{array}[]{cccc}g^{(0)}_{11}&0&0&g^{(0)}_{14}\\ 0&g^{(0)}_{22}&g^{(0)}_{23}&0\\ 0&g^{(0)}_{32}&-g^{(0)}_{22}&0\\ g^{(0)}_{41}&0&0&-g^{(0)}_{11}\\ \end{array}\right)+e^{\pi u}\times\left(\begin{array}[]{cccc}0&g_{34}&-g_{24}&0\\ g_{43}&0&0&-g_{13}\\ -g_{42}&0&0&g_{12}\\ 0&-g_{31}&g_{21}&0\\ \end{array}\right)+$ (91)
    $\quad+e^{-\pi u}\times\left(\begin{array}[]{cccc}0&g_{12}&g_{13}&0\\ g_{21}&0&0&g_{24}\\ g_{31}&0&0&g_{34}\\ 0&g_{42}&g_{43}&0\\ \end{array}\right)+e^{2\pi u}\times\left(\begin{array}[]{cccc}-g_{44}&0&0&g_{14}\\ 0&-g_{33}&g_{23}&0\\ 0&g_{32}&-g_{22}&0\\ g_{41}&0&0&-g_{11}\\ \end{array}\right)+$ (100)
@@
[…]
@@
  • We use the parametrization in terms of the 4 $\tau$-functions to force further relations. In particular
    $f(u)\equiv\mathcal{G}(u)\cdot\Theta\left(u-\mathrm{i}/2\right)$ (107)
    as $|u|\gg 1$, or
    $\lim_{u\rightarrow\pm\infty}\Theta(u)\equiv\lim_{u\rightarrow\pm\infty}\left(\kappa^{-1}\cdot f^{-1}(u)\cdot\mathcal{G}(u)\cdot\kappa\right)^{T}.$ (108)
    In (107) we demand that the expansion of $\Theta(u)$ satisfies
    $\Theta(u)\cdot\kappa\cdot\Theta^{T}(u+\mathrm{i})=\kappa$
    as $|u|\gg 1$. In this way we can demand further conditions, e.g.
    $g_{11}=g_{14}=g_{22}=g_{32}=g_{41}=g_{12}=g_{31}=g_{34}=g_{42}=0$ (109)
    together with $g_{ij}=g_{ij}(A,B,C)$ for some $(i,j)$, and relations between the parameters $A,B$ and $C$.
  • Finally, we impose the quadratic constraint (65). It implies
    $g_{22}^{(0)}=-\mathrm{i},\quad g_{32}^{(0)}=0$ (110)
    together with some supplemental relations $g_{ij}=g_{ij}(A,B,C)$ and between the coefficients $A,B$ and $C$.
Ultimately, we can parametrize the gluing matrix using only 4 parameters,
$\left(\mathcal{G}_{\text{Regge}}\right)_{i}^{\,\,\,j}=\left(\begin{array}[]{cccc}-\mathrm{i}\alpha&0&0&-\frac{\mathrm{i}\left(\alpha^{2}-1\right)}{\beta}\\ 0&-\mathrm{i}&\sigma&0\\ 0&0&\mathrm{i}&0\\ \mathrm{i}\beta&0&0&\mathrm{i}\alpha\\ \end{array}\right)+e^{\pi u}\times\left(\begin{array}[]{cccc}0&0&\gamma&0\\ -\frac{\beta\gamma\rho}{\alpha^{2}-1}&0&0&\gamma\frac{1-\rho}{\alpha}\\ 0&0&0&0\\ 0&0&-\frac{\beta\gamma\left(\alpha^{2}+\rho-1\right)}{\alpha\left(\alpha^{2}-1\right)}&0\\ \end{array}\right)+$ (119)
$e^{-\pi u}\times\left(\begin{array}[]{cccc}0&0&\gamma\frac{\rho-1}{\alpha}&0\\ -\frac{\beta\gamma\left(\alpha^{2}+\rho-1\right)}{\alpha\left(\alpha^{2}-1\right)}&0&0&-\gamma\\ 0&0&0&0\\ 0&0&-\frac{\beta\gamma\rho}{\alpha^{2}-1}&0\\ \end{array}\right)+\left(e^{2\pi u}+e^{-2\pi u}\right)\times\left(\begin{array}[]{cccc}0&0&0&0\\ 0&0&-\frac{\mathrm{i}\beta\gamma^{2}(\rho-1)}{\alpha\left(\alpha^{2}-1\right)}&0\\ 0&0&0&0\\ 0&0&0&0\\ \end{array}\right),$ (128)
where $\rho^{2}\equiv 1-\alpha^{2}$ and $\alpha,\beta,\gamma,\sigma\in\mathbb{C}$.
In the case (82) the gluing matrix solving the algorithm can be formally taken to be the transpose of the above one, with a new set of parameters, $\alpha^{\prime},\beta^{\prime},\gamma^{\prime},\sigma^{\prime}\in\mathbb{C}$,
$\left(\mathcal{G}_{\text{Bridge}}\right)_{i}^{\,\,\,j}=\left(\begin{array}[]{cccc}-\mathrm{i}\alpha^{\prime}&0&0&\mathrm{i}\beta^{\prime}\\ 0&-\mathrm{i}&0&0\\ 0&\sigma^{\prime}&\mathrm{i}&0\\ -\frac{\mathrm{i}\left(\alpha^{\prime 2}-1\right)}{\beta^{\prime}}&0&0&\mathrm{i}\alpha^{\prime}\\ \end{array}\right)+e^{\pi u}\times\left(\begin{array}[]{cccc}0&-\frac{\beta^{\prime}\gamma^{\prime}\rho^{\prime}}{\alpha^{\prime 2}-1}&0&0\\ 0&0&0&0\\ \gamma^{\prime}&0&0&-\frac{\beta^{\prime}\gamma^{\prime}\left(\alpha^{\prime 2}+\rho^{\prime}-1\right)}{\alpha^{\prime}\left(\alpha^{\prime 2}-1\right)}\\ 0&\gamma^{\prime}\frac{1-\rho^{\prime}}{\alpha^{\prime}}&0&0\\ \end{array}\right)+$ (137)
$e^{-\pi u}\times\left(\begin{array}[]{cccc}0&-\frac{\beta^{\prime}\gamma^{\prime}\left(\alpha^{\prime 2}+\rho^{\prime}-1\right)}{\alpha^{\prime}\left(\alpha^{\prime 2}-1\right)}&0&0\\ 0&0&0&0\\ \gamma^{\prime}\frac{\rho^{\prime}-1}{\alpha^{\prime}}&0&0&-\frac{\beta^{\prime}\gamma^{\prime}\rho^{\prime}}{\alpha^{\prime 2}-1}\\ 0&-\gamma^{\prime}&0&0\\ \end{array}\right)+\left(e^{2\pi u}+e^{-2\pi u}\right)\times\left(\begin{array}[]{cccc}0&0&0&0\\ 0&0&0&0\\ 0&-\frac{\mathrm{i}\beta^{\prime}\gamma^{\prime 2}(\rho^{\prime}-1)}{\alpha^{\prime}\left(\alpha^{\prime 2}-1\right)}&0&0\\ 0&0&0&0\\ \end{array}\right).$ (146)
Let us notice that the gluing matrices are defined except for a sign. Moreover, it turns out that the condition (78) is updated to

### v3

@@ 2.1.4 Asymptotics and quantum numbers
$\mathbf{P}_{ab}\simeq\#\,u^{\mathcal{N}_{a}+\mathcal{N}_{b}}\quad\text{as}\quad u\gg 1,$ (30)
@@
${\bf Q}_{ij}\simeq\#\,u^{\hat{\mathcal{N}}_{i}+\hat{\mathcal{N}}_{j}-1}\quad\text{as}\quad u\gg 1,$ (32)
@@
$\hat{M}_{1}=\Delta_{\text{top}}+S_{\text{top}}+2,\quad\hat{M}_{2}=\Delta_{\text{top}}-S_{\text{top}}+1,$
where $r_{1}\equiv 1/2\,(p_{1}+p_{2})+q$, $r_{2}\equiv 1/2\,(p_{1}+p_{2})$, $r_{3}\equiv 1/2\,(p_{2}-p_{1})$ are the corresponding $SO(6)$ weights, and in fact the condition (34) translates into the unitarity bound satisfied by physical multiplets
@@
We notice that it is the bottom component that gives the leading Regge trajectory in the sector we will consider. In the following, we will always plot only one representative for each multiplet, specifying which.
@@ 2.2 Weyl symmetries in the QSC
They appear as a symmetry of the QSC under simple Q-system transformations that multiply the “fermionic” indices of Q-functions, namely the lower indices of type $i,j=1,\dots,4$, by multiplication with one of the following matrices (each one defining a new solution):
$\left(\mathbb{M}_{1}\right)_{i}^{\,\,\,j}=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&0&1&0\\ 0&-1&0&0\\ 0&0&0&1\end{array}\right)\quad\text{or}\quad\left(\mathbb{M}_{2}\right)_{i}^{\,\,\,j}=\left(\begin{array}[]{cccc}0&0&0&1\\ 0&1&0&0\\ 0&0&1&0\\ -1&0&0&0\end{array}\right).$ (49)
In particular, the action on the Q-functions $\mathbf{Q}_{ij}$ is explicitly described by the following transformation,
$\mathbf{Q}\rightarrow\mathbb{M}\cdot\mathbf{Q}\cdot\mathbb{M}^{T},$ (50)
while upper indices, e.g. the ones in $\mathbf{Q}^{ij}$, would instead be transformed by multiplying by the inverse transposed matrices.
From the QQ-relations point of view, these transformations preserve the form of the Q-system [^14] In particular, this follows from the fact that $\mathbb{M}\cdot\kappa\cdot\mathbb{M}^{T}=\kappa$., and at the same time preserve the gluing equations (17) upon an appropriate redefinition of the gluing matrix $\mathcal{G}\rightarrow\mathbb{M}\cdot\mathcal{G}\cdot\mathbb{M}^{-1}$. However, the transformations of the form (50) also modify the asymptotics of the Q-functions, and thus allow to construct a new solution.
In particular, notice that conjugation by $\mathbb{M}_{1}$ corresponds to swapping the indices [^15] A part for the appropriate ±$\pm 1$ signs coming from the components of $\mathbb{M}$. These signs are needed to preserve the QQ-relations, but play no role in the present argument on the redefinition of the charges.
$i=2\,\leftrightarrow\,i=3,$ (51)
and thus the charges of the new solution are obtained from the original ones by the swap (the notation will be clear in a minute):
$\texttt{W}_{S}:\quad\hat{M}_{1}\,\leftrightarrow\,\hat{M}_{2}.$ (52)
@@
$\texttt{W}_{\Delta}:\quad\hat{M}_{1}\,\leftrightarrow\,-\hat{M}_{2}.$ (54)
@@
$\texttt{W}_{\Delta}:\quad\left(\Delta_{\text{bottom}},S_{\text{bottom}}\right)\,\rightarrow\,\left(3-\Delta_{\text{bottom}},S_{\text{bottom}}\right),$ (55)
@@
$\texttt{W}_{S}:\quad\left(\Delta_{\text{top}},S_{\text{top}}\right)\,\rightarrow\,\left(\Delta_{\text{top}},-1-S_{\text{top}}\right).$ (56)
These are individually Weyl reflections that leave the conformal algebra invariant. [^16] In general dimension, the two Weyl reflections would act us $\texttt{W}_{\Delta}:\,\Delta\,\leftrightarrow\,d-\Delta$, and $\texttt{W}_{S}:\,S\,\leftrightarrow\,2-d-S$. Following the CFT literature, we will call the two transformations the shadow transform and the spin-shadow. The reason why one acts on the bottom component and the other on the top component is due to the supersymmetric structure. Indeed, these two particular transformations, acting at the levels of the multiplets specified in (55) and (56), are Weyl reflections of the whole $OSp(6|4)$ superalgebra, as can be seen from the form of the Casimir eigenvalue, see e.g. Binder:2021cnk .
@@
$\tau=\left.\Delta-S\right|_{\text{top}}=\Delta_{\text{top}}-S_{\text{top}},\quad\tilde{\tau}=\left.\Delta-(\texttt{W}_{S}[S])\right|_{\text{top}}=\Delta_{\text{top}}+1+S_{\text{top}}.$ (57)
@@ 3.1 A symmetric sector
This symmetry then percolates to other elements of the construction. In particular, by studying (7), we realize that now $Q^{a|k}\kappa_{ki}$ and $\kappa^{ab}Q_{b|i}$ satisfy the same difference equation. [^17] In this paragraph, we omit $\downarrow$/$\uparrow$ / ↑ for simplicity. Identifying solutions via their asymptotics, one can choose an appropriate normalization of the Q functions such that (see, Bombardelli:2017vhk ):  [^18] We are not adopting exactly the same conventions as in Bombardelli:2017vhk . The canonical normalization of Q-functions used here for the LR-symmetric case differs from that paper. In the LR symmetric case we see that the constant $e^{i\mathcal{P}}$ defined in (29) reduces to two possible values $e^{\mathrm{i}\mathcal{P}}=\mathbf{s}=\pm 1$. A redefinition related to this sign connects the convention of this paper to the one in Bombardelli:2017vhk , i.e. $Q_{a|i}^{\text{here}}$ is equivalent to $Q_{a|i}$ in Bombardelli:2017vhk if $\mathbf{s}=1$, while if $\mathbf{s}=-1$ we multiply the $a$ index by a diagonal matrix $\left\{\mathrm{i},-\mathrm{i},-\mathrm{i},\mathrm{i}\right\}$.
@@
A second $\mathbb{Z}_{2}$ symmetry of the system is an internal symmetry acting on the spin chain representation of the states (based on the planar gauge theory diagrams) by a reflection across the midpoint of the spin chain. We will restrict our analysis to states that are eigenvalues of this transformation. At the level of Q functions this $\mathbb{Z}_{2}$ parity is realized very simply as the change of spectral parameter $u\leftrightarrow-u$ of the $\mathbf{P}$ functions. [^19] The transformation $u\leftrightarrow-u$ should be performed on a sheet where the branch cut on the real axis is drawn as short, i.e. along $[-2h,2h]$.
For parity-symmetric states, this sends the QSC solution to itself which means that the set of $\mathbf{P}$ functions is invariant under $u\leftrightarrow-u$ up to rescalings. [^20] Notice also that this symmetry implies that for parity-symmetric states the large-$u$ expansions of all quantities will go in steps of $1/u^{2}$ only. Concretely, as we review in 4.2, these functions are defined as series expansions in $1/x(u)$, where the Zhukovsky map is defined as follows:
$x(u)=\frac{u+\sqrt{u-2h}\sqrt{u+2h}}{2h}.$ (66)
In particular, $x(-u)=-x(u)$, which implies that under this parity symmetry the $\mathbf{P}$ functions pick a sign based on the charges in their asymptotics. For example, $\mathbf{P}_{12}(-u)=(-1)^{M_{1}}\mathbf{P}_{12}(u)$ since the asymptotic behavior is $\mathbf{P}_{12}(u)\simeq u^{-M_{1}}$ with $M_{1}\in\mathbb{Z}$.
We can represent these changes of sign conveniently as a Q-system transformation, i.e.
@@
Notice $\mathbb{R}\cdot\kappa\cdot\mathbb{R}^{T}=\kappa$, which ensures that this symmetry is compatible with LR symmetry.
In presence of spin chain parity, it is very simple to construct the Q-functions satisfying the Q-system holomorphic in the lower half plane, i.e. the objects we already encountered, $Q_{a|i}^{\uparrow}$ and $\mathbf{Q}_{ij}^{\uparrow}$. In fact, given the symmetry they can simply be taken as
$Q_{a|i}^{\uparrow}(u)=(\mathbb{R}^{-1})_{a}^{\,c}\,Q_{c|i}^{\downarrow}(-u),$ (69)
@@
When classifying the admissible gluing matrices compatible with this $u\leftrightarrow-u$ symmetry, we find solutions only with
$(-1)^{M_{1}+M_{2}}=-1.$ (71)
For local operators (or constant gluing matrix), this condition on the R-charges is equivalent to the spin being integer rather than half-integer (see the general quantization condition for the spins in Bombardelli:2017vhk ). In this paper we restrict our attention to trajectories interpolating between integer spins, and we leave for future work the study of more general cases.
@@ 3.1.1 Constraints on the gluing matrix
Let us now see what additional conditions should be satisfied by the gluing matrix in presence of the two symmetries above. We start by rewriting the gluing equation (17), using the $u\leftrightarrow-u$ symmetry (70), as
@@
  • Two new constraints for the matrix $\Theta(u)$ in this sector (see Appendix C), namely
@@
When imposing all constraints with these symmetry assumptions, we will find that only solutions with $(-1)^{M_{1}+M_{2}}=-1$ are allowed.
@@ 3.2.1 Strategy of the derivation
Since $\mathcal{G}(u)$ is a periodic matrix of holomorphic functions with no singularities anywhere for finite $u$, it is severely constrained. In particular, if we require that it has non-growing asymptotics at $u\rightarrow\pm\infty$, then it should be a constant by Liouville’s theorem: this is the case for QSC solutions corresponding to local operators. To have a nontrivial dependence on $u$, the most general class of functions which do not grow worse than exponentials are polynomials in $e^{\pm\pi u}$. We will make the following
@@
    $f(u)\equiv\mathcal{G}(u)\cdot\Theta\left(u-\frac{\mathrm{i}}{2}\right),$
    and impose that it can be parametrized in terms of 4 $\tau$-functions in the correct way. This equation should be studied asymptotically at both +$\infty$+ ∞ and $\infty$- ∞ .
  • Once the parameter space is sufficiently tiny, impose the quadratic constraint (76).
@@ 3.2.2 Review of the result for local operators
$\tau_{2,3}\simeq 0,$ (86)
$\tau_{1,4}\simeq\texttt{const}\,,$ (87)
@@
Using the strategy described above and in particular the algorithm explained in Appendix D, one gets as a result a matrix depending on the charges of the state (here given in $sl(2)$-like grading) and just one unfixed parameter Bombardelli:2017vhk
$\left(\mathcal{G}_{\text{local}}\right)_{i}^{\,\,\,j}=\begin{pmatrix}-\mathrm{i}\sec\left(\pi\Delta_{sl(2)}\right)&0&0&\mathrm{i}\,\delta_{1}\\ 0&\pm\,\mathrm{i}&0&0\\ 0&0&\mp\,\mathrm{i}&0\\ -\mathrm{i}\,\delta_{2}&0&0&\mathrm{i}\sec\left(\pi\Delta_{sl(2)}\right)\\ \end{pmatrix},$ (88)
@@
$\text{e}^{\imath\pi S_{sl(2)}}=\pm 1,$ (89)
where this is the same $\pm$ sign appearing in (88). We see that the cases of even/odd-spin operators are explicitly disjoint. Moreover, the following relation is required for consistency,
$(-1)^{M_{1}+M_{2}}=-1,$ (90)
in accordance with the spin being integer rather than half-integer Bombardelli:2017vhk . We note that, while here we obtained (88) under special symmetry assumptions, this form for a constant gluing matrix is actually not relegated to the particular sector we considered but is valid more generally Bombardelli:2017vhk .
@@ 3.2.3 Result for non-local operators
To avoid spin quantization, we need to relax the asymptotic behaviour of $\tau$-functions (86)-(87), which will open up the possibility for the gluing matrix to depend non-trivially on $u$. Following the strategy described above, we focus on the simplest non-trivial assumption compatible with the enhanced periodicity of the sector (64), i.e., as $u\rightarrow\pm\infty$, we allow $\tau_{2}$, $\tau_{3}$ to have a nontrivial asymptotically growing behaviour, [^21] We have included in (91)-(92) the leading and the first subleading terms. It is a general result that the corrections should have this form, due to the periodicity of the $\tau$ functions.
$\tau_{2,3}\simeq A_{2,3}^{\pm}\text{e}^{\pi\,|u|}+B_{2,3}^{\pm}\text{e}^{-\pi\,|u|},$ (91)
$\tau_{1,4}\simeq C_{1,4}^{\pm}+D_{2,3}^{\pm}\text{e}^{-2\pi\,|u|},$ (92)
where $A,B,C,D\in\mathbb{C}$ are some parameters. In this work we analyze in detail two disjoint configurations of the general ansatz (91), e.g.
@@
Notice that $A_{3}^{\pm}=A_{2}^{\pm}=0$ reduces to what discussed in Section 3.2.2.
@@
Following the strategy discussed in 3.2.1, let us examine in detail the steps we used to fix the form of the $\mathcal{G}(u)$. We describe step by step the simplifications in the case of the first assumption, the “Regge” case (93).
@@
    $\left(\begin{array}[]{cccc}g^{(0)}_{11}&0&0&g^{(0)}_{14}\\ 0&g^{(0)}_{22}&g^{(0)}_{23}&0\\ 0&g^{(0)}_{32}&-g^{(0)}_{22}&0\\ g^{(0)}_{41}&0&0&-g^{(0)}_{11}\\ \end{array}\right)+e^{\pi u}\times\left(\begin{array}[]{cccc}0&g_{34}&-g_{24}&0\\ g_{43}&0&0&-g_{13}\\ -g_{42}&0&0&g_{12}\\ 0&-g_{31}&g_{21}&0\\ \end{array}\right)$ (103)
    $\quad+e^{-\pi u}\times\left(\begin{array}[]{cccc}0&g_{12}&g_{13}&0\\ g_{21}&0&0&g_{24}\\ g_{31}&0&0&g_{34}\\ 0&g_{42}&g_{43}&0\\ \end{array}\right)+e^{2\pi u}\times\left(\begin{array}[]{cccc}-g_{44}&0&0&g_{14}\\ 0&-g_{33}&g_{23}&0\\ 0&g_{32}&-g_{22}&0\\ g_{41}&0&0&-g_{11}\\ \end{array}\right)$ (112)
@@
  • Studying equation (80) at large $u\rightarrow\pm\infty$, and exploiting the knowledge of the limit values $\Theta(\pm\infty)$ in terms of charges, we get
@@
    Up to this point, we have not yet used the specific assumptions differentiating the “Regge” and “Bridge” case.
  • We use the parametrization in terms of the 4 $\tau$-functions, (91), (92) and the “Regge” case (93), to force further relations. In particular, we impose the relations (78) and (79) at $|u|\gg 1$, which give
    $f(u)\simeq\mathcal{G}(u)\cdot\Theta\left(u-\mathrm{i}/2\right),\quad u\rightarrow\pm\infty,$ (119)
    $\lim_{u\rightarrow\pm\infty}\Theta(u+\mathrm{i}/2)\equiv\lim_{u\rightarrow\pm\infty}\left(\kappa^{-1}\cdot f^{-1}(u)\cdot\mathcal{G}(u)\kappa\right)^{T}.$ (120)